Additional baking jacket sections for the turbo section arrived today. Baking was halted at ~11:00 to allow for the chamber to cool to install new sections. New sections installed and baking restarted at 15:25. Jacket fits much better with gaps near the turbos and the motors now covered much better. Attached are some photos.
While heaters were off, the pressure fell to ~3E-9 mbar and was continuing to fall before baking was restarted.
Current baking temperature is ~75C and code is in automatic mode.
The NEG and motor controllers have been installed on the rack next to the chamber. See attached. The bottom NEG controller is connected to the 4 D-2000 NEG-ion pumps. The cables for the D-2000s have kapton tape attached. The middle NEG controller is connected to the NEG pumps whose feedthroughs are on the chamber on the inside of the ring. One of the connections on the middle controller is for the 3 NEGs connected together oppose to the usual 2. This connection is marked by white tape. The top controller is connected to the feedthrough who are on the side of the chamber on the outside of the ring. |